Find the dimension of a rectangle. The area is 238cm^ and the perimeter is 62cm?
Answer by algebrahouse.com(1659)

2x + 2y = 62 {perimeter is 2 times width + 2 times length}
xy = 238 {area is width times length}

x + y = 31 {divided each side of top equation by 2}
x = -y + 31 {subtracted y from each side}

y(-y + 31) = 238 {substituted -y + 31, in for x, into x(y) = 238}
-y² + 31y = 238 {used distributive property}
y² - 31y + 238 = 0 {added y² and subtracted 31y from each side}
(y - 17)(y - 14) = 0 {factored into two binomials}
y - 17 = 0 or y - 14 = 0 {set each factor equal to 0}
y = 17 or y = 14 {solved each equation for x}

width is 14 cm and length is 17 cm

For more help from me, visit: www.algebrahouse.com